How long will it take a 3000-W motor to lift a 405-kg piano to a sixth-story window 19.0 m above?

Express your answer to three significant figures and include the appropriate units.
Answer #1

Power of motor = 3000 W,

Mass of piano = 405 kg

Weight of piano = m×g = 405 × 9.81

Height of window = 19 m

Power = work done / time = mgh / time

Time = mgh / Power =( 405 ×9.81 × 10) / 3000

Time = 13.2435 =13.244 second